OpenStudy (anonymous):

ok i need help with a word problem. it says: 3 pails and 6 baskets contain 576 eggs.all the pails contain 1/3 as many eggs as all the baskets. how many eggs are there in each pail? how many in each basket? i have to show my work.

OpenStudy (anonymous):

Let p be the number of pails, and b the number of baskets. Now, you know that 3 pails and 6 baskets both contain 576 eggs so you form this equation. 3p + 6b = 576 We know that the pails contain a third of the eggs that the baskets contain, so we create this equation, p = (1/3)b Now, lets just replace the p from this equation to the p on the first equation and you can solve the system.

OpenStudy (anonymous):

i got that p=24 and that b=72 is that right?

OpenStudy (anonymous):

Wait, I think I have made a small mistake. The second equation is 3p=(1/3)*6b Which will give you p=48 and b=72. But good job. That was a swift solution.

OpenStudy (anonymous):

So now to answer your question you say, There are 6*b = 6*72 = 432 eggs in the baskets. There are 3*p = 3*48 = 144 eggs in the pails. In total there are 432+144 = 572 eggs.

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (anonymous):

You're welcome lynn.